# M. C. Rodriguez

### v5

@@ 4 Case 2
We can remove this Majoron, in similar way as done in our Eq.(51), when we impose that the parameters $f_{2}$, $f_{4}$, $f^{\prime}_{2}$ and $f^{\prime}_{4}$ are different from zero. In other words, we see that these terms allow us to satisfactorily implement scheme of Ma, see our Eq.(51), in MSUSY331 in a similar way as implemented in SUSYGR, see our Eqs(LABEL:suppotSUSYGR,147).
We will brief present our results when $v_{\chi}=1000$ GeV and $v_{\chi^{\prime}}=1$ GeV and we also use the following numerical parameters
@@
We also get the results shown in our Tabs.(7,8), where we notice the masses for $M_{h^{0}}$, $M_{H^{0}}$, $M_{A^{0}}$, $M_{h^{\pm}}$ and $M_{h^{\pm\pm}}$ are in agreement with the actual experimental values and we can conclude
@@ 5 Conclusions
We have studied the scalar potential of supersymmetric $331$ model with all the VEV from the sextet and anti-sextet are not null and we show our lightest scalar fields has mass near the value $M_{h^{0}}\approx 125.5$ GeV, with concordance with the experimental values for Higgs masses given by Eq.(20). We also show that in MSUSY we have $M_{A^{0}},M_{H^{0}}>400\,\ GeV$, $M_{h^{\pm}}>181$ GeV and $M_{h^{\pm\pm}}>445$ GeV. Regarding which is heavier, $H^{0}$ or $A^{0}$, the answer depends on the condition. We show that in case 1 we have $M_{A^{0}}>M_{H^{0}}$, see our Fig.(13), and in case 2 the opposite, that is, $M_{H^{0}}>M_{A^{0}}$ as we shown in our Tab.(8).
@@ Acknowledgments
Appendix A Review Scalar Potential at m331
@@
Appendix B Construction Scalar Potential

### v6

@@ 4 Case 2
We can remove this Majoron, in similar way as done in our Eq.(68), when we impose that the parameters $f_{2}$, $f_{4}$, $f^{\prime}_{2}$ and $f^{\prime}_{4}$ are different from zero. In other words, we see that these terms allow us to satisfactorily implement scheme of Ma, see our Eq.(68), in MSUSY331 in a similar way as implemented in SUSYGR, see our Eqs(LABEL:suppotSUSYGR,160).
We will brief present our results when $v_{\chi}=1000$ GeV and $v_{\chi^{\prime}}=1$ GeV and we also use the following numerical, in GeV, parameters
@@
We also get the results show in our Tabs.(6,7), where we notice the masses for $M_{h^{0}}$, $M_{H^{0}}$, $M_{A^{0}}$, $M_{h^{\pm}}$ and $M_{h^{\pm\pm}}$ are in agreement with the actual experimental values and we can conclude
@@ 5 Conclusions
We have studied the scalar potential of supersymmetric $331$ model with all the VEV from the sextet and anti-sextet are not null and we show our lightest scalar fields has mass near the value $M_{h^{0}}\approx 125.5$ GeV, with concordance with the experimental values for Higgs masses given by Eq.(28). We also show that in MSUSY331, we have $M_{A^{0}},M_{H^{0}}>400$ GeV, $M_{h^{\pm}}>181$ GeV and $M_{h^{\pm\pm}}>445$ GeV. Regarding which is heavier, $H^{0}$ or $A^{0}$, the answer depends on the condition. We show that in case 1 we have $M_{A^{0}}>M_{H^{0}}$, see our Fig.(15), and in case 2 the opposite, that is, $M_{H^{0}}>M_{A^{0}}$ as we show in our Tab.(7).
@@ Acknowledgments
Appendix A Some phenomenological Results in MSUSY331.
with
$h(t)=1+4t^{2},$ (193)
for neutrinos and
for the charged leptons, where we have defined
We can use muon collider to discover the new neutral $Z^{\prime}$ boson using the reaction $\mu e\to\mu e$ it was show at [111, 113] that $A_{RL}(\mu e)$ asymmetry is considerably enhanced, some phenomenological analyses was presented by references [101, 102, 109].
We show some values of these parameters in MSUSY331 in our Tabs.(8,9).
$v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
$v_{\chi}$ (GeV) $\xi_{WW^{\prime}}$ $\xi_{ZZ^{\prime}}$
1000 $4.09\times 10^{-4}$ $2.03\times 10^{-3}$
1500 $2.04\times 10^{-4}$ $1.25\times 10^{-3}$
2000 $1.02\times 10^{-4}$ $8.14\times 10^{-4}$
2500 $5.42\times 10^{-5}$ $5.61\times 10^{-4}$
3000 $3.04\times 10^{-5}$ $4.07\times 10^{-4}$
3500 $1.81\times 10^{-5}$ $3.07\times 10^{-4}$
4000 $1.13\times 10^{-5}$ $2.39\times 10^{-4}$
4500 $7.45\times 10^{-6}$ $1.91\times 10^{-4}$
5000 $5.06\times 10^{-6}$ $1.56\times 10^{-4}$
$g+d\rightarrow U^{--}+J,\,\ g+u\rightarrow U^{--}+j_{\alpha},$ (197)
its signature is $llXX$ and it can be detected at LHC if they really exist in nature [109, 117]. Another interesting process, with the same previous signature, to prove this model is
$e^{-}+e^{-}\rightarrow\tilde{\chi}^{--}+\tilde{\chi}^{0},$ (198)
it was presented by [^26] I want to thank M. Capdequi-Peyranère, because he taught me a lot about the phenomenology of supersymmetric models in 2000 when I started studying this model. [101, 102, 109].
Appendix B Review Scalar Potential at m331
@@
Appendix C Construction Scalar Potential